OpenStudy (anonymous):

Find the equation of the line which passes through the points (0,3), (-1, –4) in the point–slope form

OpenStudy (anonymous):

Find the slope. (y2 - y1)/(x2 - x1) = (-1-0)/(-4-3) = -1/-7 = 1/7 Point slope is: (y-y1) = m(x-x1) Pick one set of coordinates to plug in for x1 and y1, m is slope, as above. Using (0-3): y - (-3) = (1/7)(x - 0) y + 3 = (1/7)x Subtract 3 from both sides. y + 3 - 3 = (1/7)x - 3 y = (1/7)x - 3

OpenStudy (anonymous):

thanks so much! but why is -1/-7 not just -1/-7 instead of 1/7?

OpenStudy (anonymous):

I used (0, -3) instead of (0,3) so just make the substitution. -1/-7 is a negative divided by a negative, so you get a positive.